A penguin slides down a 12m slope. If it takes the penguin 4s to reach the bottom, What was the penguin's velocity?
1
Expert's answer
2011-10-14T08:50:05-0400
The motion is accelerated (by gravity force) hence the velosity v=v0+at, a -acceleration
the way travelled s=s0+v0t+at^2/2
we assume that the initial velosity of the penguin v0=0
hence s=at^2/2
a=2s/t^2=1.5m/s^2
v=1.5*4=6 m/s - the final velosity at the end of the slope
motion is accelerated so the velosity is changing during penguin's sliding